# ReID5o: Achieving Omni Multi-modal Person Re-identification in a Single Model

### v1

@@ Abstract
In real-word scenarios, person re-identification (ReID) expects to identify a person-of-interest via the descriptive query, regardless of whether the query is a single modality or a combination of multiple modalities. However, existing methods and datasets remain constrained to limited modalities, failing to meet this requirement. Therefore, we investigate a new challenging problem called Omni Multi-modal Person Re-identification (OM-ReID), which aims to achieve effective retrieval with varying multi-modal queries. To address dataset scarcity, we construct ORBench, the first high-quality multi-modal dataset comprising 1,000 unique identities across five modalities: RGB, infrared, color pencil, sketch, and textual description. This dataset also has significant superiority in terms of diversity, such as the painting perspectives and textual information. It could serve as an ideal platform for follow-up investigations in OM-ReID. Moreover, we propose ReID5o, a novel multi-modal learning framework for person ReID. It enables synergistic fusion and cross-modal alignment of arbitrary modality combinations in a single model, with a unified encoding and multi-expert routing mechanism proposed. Extensive experiments verify the advancement and practicality of our ORBench. A wide range of possible models have been evaluated and compared on it, and our proposed ReID5o model gives the best performance. The dataset and code will be made publicly available.
@@ 1 Introduction
We propose that exploring such a challenging problem, termed Omni Multi-modal Person Re-identification (OM-ReID), is of great significance. On the one hand, due to the diversity of information acquisition, there is often a need for multi-modal queries in practical applications [50]. Users expect the model to be able to effectively process multi-modal information and provide comprehensive and accurate retrieval results. If the model can handle these varying modalities, it will significantly boost its practicality. On the other hand, theoretically, different modalities provide abundant and complementary features [53, 5, 21]. For instance, RGB images are vulnerable to environmental light
@@
variations [46], and infrared as well as sketch images lack essential color information critical for ReID tasks. Fusing multi-modal features is akin to assembling an information jigsaw from diverse perspectives and dimensions. This enables a more comprehensive portrayal of person characteristics, thereby minimizing misjudgments stemming from single-modal limitations.
@@ 2.1 Construction Motivation
Multi-modal learning [52, 44, 61, 19] has become a hot topic in the computer vision field. However, when directing to person ReID, researchers often confine themselves to datasets with few modalities and unsatisfactory quality [24, 34, 53, 5, 21]. Therefore, we construct ORBench, the first high-quality dataset with five modalities, driven by the following two major motivations.
@@ 2.4 Evaluation Protocol
As a common practice, we adopt Cumulative Matching Characteristic (CMC) [32] and mean average precision (mAP) to evaluate the performance. Given a query, all gallery images are ranked according to their affinities with the query. A successful search is achieved if any image of the corresponding person is among the rank-k images.
@@ 3 ReID5o: Method
3.1 Overview
In this section, we propose a unified multi-modal learning framework for person ReID, named ReID5o. We utilize a multi-modal tokenizing assembler to encode the multi-modal inputs into a shared embedding space. Then, a unified encoder is used to extract modality-shared features, and a multi-expert router is designed to facilitate modality-specific representation learning. Simultaneously, a feature mixture is employed to fuse the multi-modal features. Finally, we utilize a simple strategy to implement cross-modal alignment, using the SDM [14] and identity classification losses. The whole schematic is shown in Fig. 3.
@@ 3.3 Multi-Expert Router
We utilize a unified multi-modal encoder to extract modality-shared features, which inherits the rich multi-modal pre-trained knowledge from CLIP [38] to ensure a good initial starting point for subsequent training. As a Transformer, it contains multiple linear layers. For a certain linear layer $W\in\mathbb{R}^{d\times k}$, its function can be reduced as $\textbf{y}=W(\textbf{x})$, where $\textbf{x}\in\mathbb{R}^{d}$ and $\textbf{y}\in\mathbb{R}^{k}$ represent the input and output feature for this layer, respectively. It can be seen that if the same linear layer is directly used to extract features for different modalities, it will be hard to extract the modality-specific representations.
@@ 3.4 Feature Mixture and Learning Strategy
Now, for the input data $\mathbf{X}^{mod}$, through the aforementioned tokenization and expert routing mechanisms, we can obtain its output features $\mathbf{Z}^{mod}\in\mathbb{R}^{n\times D}$. Then, to achieve full complementary information interaction between different modalities, we employ an efficient feature mixture to fuse the features of varying modality combinations. The feature mixture consists of a multi-head self attention (MSA) layer, 1-layer transformer block and a multi-layer perception (MLP). For each multi-modal features $\mathbf{Z}^{mod}$, we perform a sequential combinatorial traversal. After that, we concatenate the combined features and feed them into our feature mixer. In this way, we can obtain the fused features.
@@ 4.1 Experimental Settings
Implementation Details. We employ a pre-trained multi-modal encoder [38], i.e., CLIP-B/16, as the unified encoder, which is pre-trained on billions of image-text pairs with contrastive learning. For the multi-modal tokenizing assembler (MTA), the tokenizers for all modalities are initialized simultaneously with the pre-trained parameters in CLIP [38], and will not be shared during the subsequent training. For the multi-expert router (MER), each modality will be assigned an independent expert, and the low-rank matrices [13] of a certain expert will be installed on each linear layer of the encoder, where $r$ is set to 4. For each layer of the feature mixture (FM), the hidden size and number of heads are set to 512 and 8. The hyper-parameter $\alpha$ for the ID loss is set to 1.0. During training, the images of different modalities are uniformly resized to $384\times 128$ and the maximum length of the textual tokens is set to 77. Also, random erasing, horizontally flipping and crop with padding are employed for image augmentation. Random masking and replacement is employed for text augmentation. Our ReID5o is trained with Adam [17] for 60 epochs with an initial learning rate $1e^{-5}$. We spend 5 warm-up epochs linearly increasing the learning rate from $1e^{-6}$ to $1e^{-5}$. For the random-initialized experts and feature mixture, the initial learning rate is set to $5e^{-5}$. The experiments are performed on a single A100 80GB GPU.
@@ 5 Conclusion
This paper tackles the challenging task of OM-ReID, focusing on cross-modality retrieval across diverse data types. We introduce ORBench, the first high-quality multi-modal dataset for this task, comprising 5 modalities for 1,000 identities. Unlike prior datasets, ORBench captures comprehensive and diverse individual portrayals, enabling research on arbitrary modality combinations. We also propose ReID5o, a novel multi-modal framework designed to handle flexible modality inputs, allowing effective retrieval across single or combined modalities. Experiments show that multi-modal combination querying significantly outperforms single-modality approaches in real-world scenarios. By providing ORBench and ReID5o, this work establishes a foundational baseline for OM-ReID research. We hope our dataset and framework will inspire further advancements in person ReID.
@@ Appendix A Appendix
A.1 Related Work
Cross-modal Person ReID. Compared to single-modal person ReID tasks [20, 30, 56], cross-modal ReID tasks [5, 46] typically use RGB images as retrieval targets while employing text descriptions [25, 39, 9, 48, 26], infrared images [37, 49, 47], attributes [41], or sketches [11, 6] as queries to simulate real-world scenarios with uncertain query modalities. Li et al.[25] pioneered text-based person ReID. Wang et al.[41] proposed a joint attribute-identity method to learn attribute-semantic and identity-discriminative features. Qiu et al.[37] introduced a high-order structure-based intermediate feature learning network to fully exploit structural information in IR features. Lin et al.[29] designed mechanisms to mitigate the impact of sketches painter’s subjectivity on ReID performance. Subsequent work combines some modalities to address practical constraints in obtaining target modalities and using complementary information. Chen et al.[5] fused features of RGB, sketch, and text modalities for modality-agnostic retrieval. Li et al.[22] developed a unified framework handling RGB, infrared, sketch, and text modalities. However, existing works are still constraint to scenarios with few modalities and overlook the rich cross-domain information contained in different modality combinations.
A.2 More Specific Samples
@@
A.3 Public Evaluation Procedure
@@
A.4 More Experimental Results
@@
A.5 Broad Impact Discussion
This paper explores the multi-modal person re-identification technology. Our research can help to identify and track persons across different cameras, thus boosting the development of smart retail, smart transportation, smart security systems and so on in the future metropolises.
Nevertheless, the application of person re-identification, such as for identifying and tracking persons in surveillance systems, might raise privacy concerns. It typically depends on the utilization of surveillance data for training without the explicit consent of the individuals being recorded. Therefore, governments and officials need to carefully establish strict regulations and laws to control the usage of this technology. Otherwise, this technology can potentially equip malicious actors with the ability to surveil pedestrians through multiple cameras without their consent. Meanwhile, the research community should also avoid using any dataset with ethics issues. For example, DukeMTMC has been taken down due to the violation of data collection terms and should no longer be used. Furthermore, we should be cautious of the misidentification of such person retrieval systems to avoid possible disturbance. Also, note that the demographic makeup of the datasets used is not representative of the broader population.
At the same time, we have utilized some public datasets containing persons for constructing our dataset and conducting the experiments. Consequently, it is inevitable that the resulting sources may inherently contain information about these persons. Researchers should adhere to relevant laws and regulations, and strive to avoid using our sources for any improper invasion of privacy. We will have a gated release of our dataset and code to avoid any misuse. We will require that users adhere to usage guidelines and restrictions to access our sources. Meanwhile, all our open-sourced sources can only be used for research purpose and are forbidden for any commercial use.
A.6 Controlled Release
@@
A.7 Limitation
Although our ReID5o demonstrates superior multi-modal retrieval capabilities, particularly on our constructed ORBench dataset, we acknowledge that our work still has some limitations. On the one hand, in real-world application scenarios, color painting and sketch queries used to search for targets often contain significant noise—that is, these artistic-style queries may not be highly consistent with the person being searched. Our ORBench ensures high data quality (identity consistency) through strict manual screening, which to some extent does not account for noise in real-world scenarios and makes it difficult to evaluate the model’s true performance in noisy environments. Additionally, there are some modalities that are not fully included in our dataset, such as thermal infrared modality, radar modality, and others. In our future work, we will focus on researching how to measure and enhance the model’s capabilities in real-world noisy multi-modal retrieval scenarios. Nevertheless, we firmly believe that our current ORBench dataset and ReID5o model can bring substantial inspiration and contributions to the field of person ReID.

### v2

# Re-identification in a Single Model

@@ Abstract
In real-word scenarios, person re-identification (ReID) expects to identify a person-of-interest via the descriptive query, regardless of whether the query is a single modality or a combination of multiple modalities. However, existing methods and datasets remain constrained to limited modalities, failing to meet this requirement. Therefore, we investigate a new challenging problem called Omni Multi-modal Person Re-identification (OM-ReID), which aims to achieve effective retrieval with varying multi-modal queries. To address dataset scarcity, we construct ORBench, the first high-quality multi-modal dataset comprising 1,000 unique identities across five modalities: RGB, infrared, color pencil, sketch, and textual description. This dataset also has significant superiority in terms of diversity, such as the painting perspectives and textual information. It could serve as an ideal platform for follow-up investigations in OM-ReID. Moreover, we propose ReID5o, a novel multi-modal learning framework for person ReID. It enables synergistic fusion and cross-modal alignment of arbitrary modality combinations in a single model, with a unified encoding and multi-expert routing mechanism proposed. Extensive experiments verify the advancement and practicality of our ORBench. A range of models have been compared on it, and our proposed ReID5o gives the best performance.
@@ 1 Introduction
We propose that exploring such a challenging problem, termed Omni Multi-modal Person Re-identification (OM-ReID), is of great significance. On the one hand, due to the diversity of information acquisition, there is often a need for multi-modal queries in practical applications [52]. Users expect the model to be able to effectively process multi-modal information and provide comprehensive and accurate retrieval results. If the model can handle these varying modalities, it will significantly boost its practicality. On the other hand, theoretically, different modalities provide abundant and complementary features [54, 5, 24]. For instance, RGB images are vulnerable to environmental light variations [48], and infrared as well as sketch images lack essential color information critical for ReID tasks. Fusing multi-modal features is akin to assembling an information jigsaw from diverse perspectives and dimensions. This enables a more comprehensive portrayal of person characteristics, thereby minimizing misjudgments stemming from single-modal limitations.
@@ 2.1 Construction Motivation
Multi-modal learning [46, 62, 22] has become a hot topic in the computer vision field. However, when directing to person ReID, researchers often confine themselves to datasets with few modalities and unsatisfactory quality [27, 36, 54, 5, 24]. Therefore, we construct ORBench, the first high-quality dataset with five modalities, driven by the following two major motivations.
@@ 2.4 Evaluation Protocol
As a common practice, we adopt Cumulative Matching Characteristic (CMC) and mean average precision (mAP) to evaluate the performance. Given a query, all gallery images are ranked according to their affinities with the query. A successful search is achieved if any image of the corresponding person is among the rank-k images.
2.5 Discussion on Realism and Idealized Assumptions.
It is important to acknowledge the idealized assumptions underpinning our dataset construction. We operated under the premise that comprehensive eyewitness information enables the creation of a complete textual description and, subsequently, a high-fidelity visual rendering of the target. Consequently, the identity consistency across modalities in ORBench is intentionally high.
In the taxonomy of [20], our sketches and color paintings are best characterized as "viewed sketches"—precise depictions derived from ample information—rather than incomplete or inaccurate "forensic sketches" that result from the fallibility of human memory. This design choice was necessitated by the feasibility of large-scale dataset construction. While ORBench provides a foundational, high-quality benchmark for studying multi-modal alignment in a controlled setting, it does not fully capture the noise and variance inherent in real-world forensic scenarios. We posit that this clean dataset serves as a crucial starting point and a powerful pre-training resource.
@@ 3 ReID5o: Method
In this section, we propose a unified multi-modal learning framework for person ReID, named ReID5o. We utilize a multi-modal tokenizing assembler to encode the multi-modal inputs into a shared embedding space. Then, a unified encoder is used to extract modality-shared features, and a multi-expert router is designed to facilitate modality-specific representation learning. Also, a feature mixture fuses the multi-modal features. Finally, we use the SDM [16] and identity classification losses to implement cross-modal alignment. The whole schematic is shown in Fig. 3.
@@ 3.2 Multi-Expert Router
We utilize a unified multi-modal encoder to extract modality-shared features, which inherits the rich multi-modal pre-trained knowledge from CLIP [40] to ensure a good initial starting point for subsequent training. As a Transformer, it contains multiple linear layers. For a certain linear layer $W\in\mathbb{R}^{d\times k}$, its function can be reduced as $\textbf{y}=W(\textbf{x})$, where $\textbf{x}\in\mathbb{R}^{d}$ and $\textbf{y}\in\mathbb{R}^{k}$ represent the input and output feature for this layer, respectively. It can be seen that if the same linear layer is used directly for different modalities, it is hard to extract modality-specific representations.
@@ 3.3 Feature Mixture and Learning Strategy
Now, for the input data $\mathbf{X}^{mod}$, through the aforementioned tokenization and expert routing mechanisms, we can obtain its output features $\mathbf{Z}^{mod}\in\mathbb{R}^{n\times D}$. Then, to achieve full complementary information interaction between different modalities, we employ an efficient feature mixture to fuse the features of varying modality combinations. The feature mixture consists of a multi-head self attention (MSA) layer, 1-layer transformer block and a multi-layer perception (MLP). Then, the combinatorial traversal is performed on the entire set of multimodal query features $\{Z^{I},Z^{C},Z^{S},Z^{T}\}$. We generate all possible combinations from single-modal to quad-modal. These combinations are then concatenated and processed by the Feature Mixture module.
@@ 4.1 Experimental Settings
Implementation Details. We employ a pre-trained multi-modal encoder [40], i.e., CLIP-B/16, as the unified encoder, which is pre-trained on billions of image-text pairs with contrastive learning. In the actual implementation, to fully leverage pre-trained knowledge, we employ an independent text encoder for the textual modality, while adopting the aforementioned strategy for other visual modalities. For the multi-modal tokenizing assembler (MTA), the tokenizers for all modalities are initialized simultaneously with the pre-trained parameters in CLIP [40], and will not be shared during the subsequent training. For the multi-expert router (MER), each modality will be assigned an independent expert, and the low-rank matrices [15] of a certain expert will be installed on each linear layer of the encoder, where $r$ is set to 4. For each layer of the feature mixture (FM), the hidden size and number of heads are set to 512 and 8. The hyper-parameter $\alpha$ for the ID loss is set to 1.0. During training, all images are uniformly resized to $384\times 128$ and the maximum length of the textual tokens is set to 77. Also, random erasing, horizontally flipping and crop with padding are employed for image augmentation. Random masking and replacement is employed for text augmentation. Our ReID5o is trained with Adam [19] for 60 epochs with an initial learning rate $1e^{-5}$. We spend 5 warm-up epochs linearly increasing the learning rate from $1e^{-6}$ to $1e^{-5}$. For the random-initialized experts and feature mixture, the initial learning rate is set to $5e^{-5}$. We use a single A100 80GB GPU.
@@ 4.6 Visualization
5 Related Work
Cross-modal Person ReID. Compared to single-modal person ReID tasks [23, 33, 57, 7, 14], cross-modal ReID tasks [5, 48] typically use RGB images as retrieval targets while employing text descriptions [28, 41, 10, 50, 29], infrared images [39, 51, 49], attributes [43], or sketches [12, 6] as queries to simulate real-world scenarios with uncertain query modalities. Li et al.[28] pioneered text-based person ReID. Wang et al.[43] proposed a joint attribute-identity method to learn attribute-semantic and identity-discriminative features. Qiu et al.[39] introduced a high-order structure-based intermediate feature learning network to fully exploit structural information in IR features. Lin et al.[32] designed mechanisms to mitigate the impact of sketches painter’s subjectivity on ReID performance. Subsequent work combines some modalities to address practical constraints in obtaining target modalities and using complementary information. Chen et al.[5] fused features of RGB, sketch, and text modalities for modality-agnostic retrieval. Li et al.[25] developed a unified framework handling RGB, infrared, sketch, and text modalities. However, existing works are still constraint to scenarios with few modalities and overlook the rich cross-domain information contained in different modality combinations.
6 Limitation
While the proposed ReID5o framework and the ORBench dataset establish a strong foundation for OM-ReID research, we acknowledge several limitations that point to future directions. First and foremost, ORBench is constructed under an idealized data generation assumption. The high fidelity of its sketch and color painting queries, which resemble "viewed sketches" rather than imperfect "forensic sketches," ensures tight cross-modal alignment but does not fully encapsulate the noise and variance (e.g., inaccuracies, omissions) prevalent in real-world scenarios. Consequently, while our model demonstrates robust performance on this clean benchmark and shows promising generalization in cross-dataset tests, its performance in truly noisy, in-the-wild settings requires further investigation. Secondly, our work is currently limited to a predefined set of five modalities. Other practically relevant modalities, such as thermal infrared or radar, are not included, and our framework is not yet designed for incremental learning of new modalities. Finally, the study focuses on a closed-set identification task, leaving open-world scenarios with unseen identities as an open challenge. We believe addressing these limitations—by introducing realistic noise, extending modality coverage, and moving towards open-set recognition—will be crucial steps for transitioning OM-ReID from a controlled benchmark to robust real-world applications.
@@ 7 Conclusion
This paper investigated the problem of Omni Multi-modal Person Re-identification (OM-ReID), where the goal is to retrieve a person using a query from any single modality or their arbitrary combinations. To support this research, we constructed ORBench, the first comprehensive dataset for OM-ReID, featuring 1,000 identities across five modalities (RGB, infrared, sketch, color pencil, and text). We also developed ReID5o, a flexible framework that enables unified encoding and effective retrieval for any modality input. Our experiments demonstrate that leveraging multi-modal queries substantially improves retrieval performance compared to single-modality approaches, underscoring their practical value. We believe that the ORBench dataset and the ReID5o framework establish a solid foundation and a strong baseline for future research in this emerging field.
8 Acknowledgments.
This work was supported by the National Natural Science Foundation of China No.62176097, and the Hubei Provincial Natural Science Foundation of China No.2022CFA055. We would like to sincerely thank the HPC Platform of Huazhong University of Science and Technology for providing computational resources.
@@ Appendix A Appendix
A.1 More Specific Samples
@@
A.2 Public Evaluation Procedure
@@
A.3 More Experimental Results
@@
A.4 Controlled Release
@@
A.5 Broad Impact Discussion
This paper explores multi-modal person re-identification technology, which holds promise for applications in smart retail, intelligent transportation, and public safety systems by enabling cross-camera identification and tracking of individuals in urban environments.
However, the deployment of such technology must be approached with careful consideration of its ethical and societal implications. First, we explicitly acknowledge that the term “bias” in this context encompasses not only labeled sensitive attributes but also imbalances in data distribution. Our dataset, ORBench, is sourced primarily from university settings and does not reflect a globally uniform demographic distribution. Certain demographic groups—such as the elderly, individuals with visible disabilities, or specific racial groups—are likely underrepresented. As a result, model performance may vary across populations, and end-users should validate the fairness and generalization of the system in their specific application contexts prior to deployment.
Furthermore, while we adopted a structured annotation protocol to maximize objectivity and consistency—focusing on descriptive visual attributes such as clothing and accessories—we recognize that the design of such guidelines inherently introduces a form of methodological bias. By prescribing which features are salient, the annotation process shapes the model’s attention and may omit other potentially relevant characteristics.
The potential for misuse of this technology also warrants serious attention. Although our research is intended for academic and public safety purposes, the capability to retrieve individuals via textual descriptions could be repurposed to target people based on attributes associated with protected categories—such as disability, religious attire, or age. Like many dual-use technologies, person re-identification systems require strong governance, ethical oversight, and clear usage policies to prevent discriminatory or privacy-infringing applications. We oppose any such misuse.
In light of these considerations, we emphasize the importance of legal and regulatory frameworks to govern the application of ReID systems. Training often relies on surveillance data collected without explicit individual consent, raising privacy concerns. The research community should avoid using ethically problematic datasets and adopt responsible data practices. To mitigate potential harm, we will release ORBench under gated access, with usage restricted to non-commercial research purposes and subject to strict guidelines.
By addressing these issues transparently, we hope to encourage the responsible development and deployment of multi-modal person re-identification systems.